# Masks and Boxes: Combining the Best of Both Worlds for Multi-Object Tracking

### v2

@@ F.2 Mask confidence
Each predicted pixel in the mask provided by the mask temporal propagator is assigned a confidence probability [8, 6]. To determine if the mask prediction is reliable enough to be used in the tracklet-detection association process, we average these confidence values across all pixels for each tracklet mask separately. If the average confidence falls below our threshold (set to 0.6, as explained in Sec. 4.1 of the main paper), the association relies solely on the intersection over union (IoU) score.
F.3 Bounding box coverage (mm1) and mask fill ratio (mm2)
@@
Figure 6: Cases showing the differences in $mm_{1}$ and $mm_{2}$ (Sec. F.3) values of a mask (in blue) within a bounding box. The most optimal case for the mask to provide a good guidance is the second one from the left, where both $mm_{1}$ and $mm_{2}$ are as close to $1$ as possible.
@@
For a more detailed explanation of $mm_{1}$ and $mm_{2}$ (see Sec. 3.3 in the main paper), we refer to the accompanying figure that illustrates their varying values. For convenience, this figure is also included in the supplementary material, see Fig. 6.
@@
In the main paper, we define the bounding box coverage of the mask, referred to as $mm_{1}$. It measures the percentage of the tracklet’s mask that falls within the detection’s bounding box for each tracklet-detection pair (if the mask is visible and confident). When the entire mask is within the detection’s bounding box, the $mm_{1}$ value is 1.0, as seen in the second and third case from the left in Fig. 6. However, if part of the mask extends outside the bounding box, the $mm_{1}$ value decreases, as shown in the first and fourth case from the left in Fig. 6. Bounding boxes can sometimes be slightly inaccurate, so we allow the $mm_{1}$ value to be slightly below 1.0, with a threshold of 0.9 (see Sec. 4.1 in the main paper). However, if too much of the mask extends beyond the bounding box, it could indicate that the mask belongs to a different detection, in which case the mask is excluded from influencing the cost of the tracklet-detection match.
@@
In the main paper, we also define the mask fill ratio of the bounding box, referred to as $mm_{2}$. This is calculated for each tracklet-detection pair and measures the percentage of the detection bounding box covered by the mask. If the mask only covers a small portion of the bounding box, the $mm_{2}$ value will be low, as illustrated in the third case from the left in Fig. 6. This can indicate a noisy or incorrect mask prediction or that part of another mask is overlapping the bounding box. To avoid misleading associations, we set a minimum threshold for $mm_{2}$ (0.05, see Sec. 4.1 in the main paper) to ensure reliable guidance.
@@
For higher coverage, the $mm_{2}$ value naturally increases. However, this value alone is not enough to confirm whether the mask tracklet properly matches the detection bounding box. For example, in the first case from the left in Fig. 6, the mask covers most of the bounding box, but it extends significantly beyond it. In the fourth case, $mm_{2}$ is 1.0, indicating the mask fully covers the bounding box, but it sticks out even more, making it an unlikely good match. Such situations can occur when one subject is covered by another, closer to the camera. While $mm_{2}$ can approach 1.0, it is rare to see exactly 1.0 in good matches, as subjects (such as people) are not rectangular and do not fully align with their bounding boxes.
@@
For the mask to have the best influence on the tracklet-detection pair match, both $mm_{1}$ and $mm_{2}$ should be as close to 1.0 as possible, as shown in the second case from the left in Fig. 6.
@@
Since high $mm_{1}$ values can be misleading when a mask significantly extends beyond the bounding box, $mm_{1}$ is used only as a gating condition to prevent poor matches (e.g., the fourth case from the left in Fig. 6). Instead, we use $mm_{2}$, which better reflects the match quality between a tracklet mask and a detection bounding box, provided $mm_{1}$ is sufficiently high.
F.4 Camera motion compensation

### v3

@@ E.3 Mask confidence
Each predicted pixel in the mask provided by the mask temporal propagator is assigned a confidence probability [6, 8]. To determine if the TP mask prediction is reliable enough to be used in the tracklet-detection association process, we average these confidence values across all pixels for each tracklet mask separately. If the average confidence is too low, the association relies solely on the intersection over union (IoU) score.
@@
Figure 7: Cases showing the differences in $mm_{1}$ and $mm_{2}$ values of a temporally propagated mask (in blue) within a bounding box (Sec. E.4). The most optimal case for the mask to provide a good guidance is the second one from the left, where both $mm_{1}$ and $mm_{2}$ are as close to $1$ as possible.
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 8: Visual output comparison between the baseline and McByte. With the temporally propagated mask guidance, McByte can handle the association of an ambiguous set of bounding boxes - see the subjects with IDs 59 and 63 on the output of McByte. Input image data from [28]. Best seen in color.
E.4 Bounding box coverage (mm1) and mask fill ratio (mm2)
@@
For a more detailed explanation of $mm_{1}$ and $mm_{2}$ (see Sec. 3.3 in the main paper), we refer to the accompanying figure that illustrates their varying values. For the reader’s convenience, this figure is also included in the supplementary material, see Fig. 7.
@@
In the main paper, we define the bounding box coverage of the mask, referred to as $mm_{1}$. It measures the percentage of the tracklet’s TP mask that falls within the detection’s bounding box for each tracklet-detection pair (if the mask is visible and confident). When the entire TP mask is within the detection’s bounding box, the $mm_{1}$ value is 1.0, as seen in the second and third case from the left in Fig. 7. However, if part of the TP mask extends outside the bounding box, the $mm_{1}$ value decreases, as shown in the first and fourth case from the left in Fig. 7. Bounding boxes can sometimes be slightly inaccurate, so we allow the $mm_{1}$ value to be slightly below 1.0. However, if too much of the TP mask extends beyond the bounding box, it could indicate that the mask belongs to a different detection, in which case the mask is excluded from influencing the cost of the particular tracklet-detection pair match.
@@
In the main paper, we also define the mask fill ratio of the bounding box, referred to as $mm_{2}$. This is calculated for each tracklet-detection pair and measures the percentage of the detection bounding box covered by the TP mask. If the TP mask only covers a small portion of the bounding box, the $mm_{2}$ value will be low, as illustrated in the third case from the left in Fig. 7. This can indicate a noisy or incorrect mask prediction or that part of another TP mask is overlapping the bounding box. In our conditions (Sec. 3.3 of the main paper) we do not allow the mask to influence the tracklet-detection association in such cases.
@@
For higher coverage, the $mm_{2}$ value naturally increases. However, this value alone is not enough to confirm whether the tracklet, to which the TP mask is linked, properly matches the detection bounding box. For example, in the first case from the left in Fig. 7, the TP mask covers most of the bounding box, but it extends significantly beyond it. In the fourth case, $mm_{2}$ is 1.0, indicating that the TP mask fully covers the bounding box, but it sticks out even more, making it an unlikely good match. Such situations can occur when one subject is covered by another, closer to the camera. While $mm_{2}$ can approach 1.0, it is rare to see exactly 1.0 in good matches, as subjects (such as people) are not rectangular and do not fully align with their bounding boxes.
@@
For the TP mask to have the best influence on the tracklet-detection pair match, both $mm_{1}$ and $mm_{2}$ should be as close to 1.0 as possible, as shown in the second case from the left in Fig. 7.
@@
Since high $mm_{1}$ values can be misleading when a TP mask significantly extends beyond the bounding box, $mm_{1}$ is used only as a gating condition to prevent poor matches (e.g., the fourth case from the left in Fig. 7). Instead, we use $mm_{2}$, which better reflects the match quality between a tracklet TP mask and a detection bounding box, provided $mm_{1}$ is sufficiently high.
E.5 Camera motion compensation
@@
E.6 Additional visual example
In the main paper, we discuss that McByte can handle challenging scenarios due to the temporally propagated mask signal used in the controlled manner as an association cue (Sec. 3.3). Fig. 8 in this supplementary material illustrates our method handling association of ambiguous boxes, improving over the baseline. A figure illustrating handling long term occlusions in the crowd is placed in the main paper (Sec. 3.3).